OpenStudy (anonymous):

charles opened the trunk and found $6750 in $1 bills and $10 bills. If there were 150 more ones than tens, how many of each kind were there?

OpenStudy (anonymous):

You need to set up a system of equations here. As you're looking for the number of each kind of bill, we'll make a variable for each kind: let \(o\) be the number of ones, and \(t\) be the number of tens. First, we know that together they make \($6750\). Second, we know how many more ones there are then tens: \[ \begin{align} o + 10t = 6750\\ o + 150 = t\\ \end{align} \] Now all you need to do is solve the system for \(o\) and \(t\).

OpenStudy (anonymous):

let the 10s be x while 1s be y 10x+y=6750 11x=6600 + x-y=150 x=600, y=450